Find the equation of the line in STANDARD FORM through the points (2,1) and (-14,5)
Answer by Alan3354(69443)

Find the equation of the line in STANDARD FORM through the points (2,1) and (-14,5)
----------------
Do it like this:
Find the equation given 2 points:
Points (3,7) and (6,1)
Find m, the slope
m = diffy/diffx = (6-3)/(1-7) = 3/(-6)
m = -1/2
--------
Use y = mx + b and either point to find b, the y-intercept
7 = (-1/2)*3 + b
b = 8.5 or 17/2
--> y = (-1/2)x + 17/2 in slope-intercept form
x + 2y = 17 in standard form